OpenStudy (anonymous):

checkiiddiie my work please? What is the simplified form of a over 3 all over 6 over 3a ? 6 over a squared 1 a squared over 6 6a I got D!

jimthompson5910 (jim_thompson5910):

So the expression is \[\large \frac{\frac{a}{3}}{\frac{6}{a}}\] right?

OpenStudy (anonymous):

yes sir

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

Im right?

jimthompson5910 (jim_thompson5910):

oh wait, i missed a 3a, one sec

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

unfortunately it is incorrect

OpenStudy (anonymous):

can you explain please?

jimthompson5910 (jim_thompson5910):

what you need to do is flip the second fraction in the bottom and multiply

jimthompson5910 (jim_thompson5910):

so  turns into